THE ROTATIONAL SPECTRA OF DIMETHYLDISELENIDE \underline{C. W. GILLIES}, Department of Chemistry, Rensselaer Polytechnic Institute, troy, NY 12180; J. Z. GILLIES and J. POLLINO, Department of Chemistry, Siena College, Loudonville, NY 12211.